### v2

@@ 4.5 Privacy Expert Chat
One participant asked if the backup should be encrypted during Task~3.2. For Task~3.2, two participants asked the expert to clarify if data from the backup also needed to be deleted. For Task~3.2, two participants asked about data minimization, purpose limitation, and access control measures for doctors. One participant contacted the expert with a technical question, as they deleted more data from the backup by accident and were unsure how to resolve the issue. Lastly, concerning Task~3.2, two participants asked questions regarding consent. One participant asked the expert to clarify if the users already had consented to advertisements and to confirm that, without consent, the sending of the emails would not be compliant:
%But if a user has not given any consent then sending mails to them would not be privacy-compliant or is it?m — E1 Chat
@@
%Would you consider it to be okay just to send all the existing users the advertisement email, but include a link to opt-out on future emails of that kind? I am personally still concerned about such an approach, but you are the expert regarding privacy topics - so I’m happy to hear your advicem — E9 Chat
@@
%I asked the one time, if all the data, if the patient’s data should also be deleted from the backups, but it was the expert, of course, said yes, and it was kind of obvious that it should be deletedm — E6
@@
%Thanks again and sorry for bugging you so much latelym — E9 Chat
@@
%It would have been easier for me if either the task was basically in line with data protection or if I had received more precise instructions from the expertm — E9
@@
%Cause I’m kind of afraid of asking too muchm — E2
@@
%That’s funny. Yeah. Maybe I was just too lazy and wanted to get the tasks donem — E4
@@
%I don’t know because I have no… Yeah, I should have used it. Yeah, I think I should have used it.m — E2
@@ 4.6 Confidence
Twenty participants found the programming tasks moderately challenging. Two (P6, E5) found them not challenging at all, and three (C1, C3, E3) found them slightly challenging. Four (C7, P5, E1, E6) found the tasks very challenging, and for one participant (C4), the tasks were extremely challenging. An overview of the self-evaluation for each participant can be found in Table V.
@@ 4.8 Participants’ Explanation for Privacy Behavior
%I just found something, but I haven’t written it because it’s already in the project, so it’s not my faultm — E2
indicating that participants while noticing privacy issues, decided not to address them as they were not seen as part of a developer’s responsibility. For example, one participant commented on Task~3.2:
%I’m not sure about whether it’s okay to send everyone an email just randomly. So I don’t know whether you had to make some subscription to something like a newsletter or something because, yeah, I don’t know. And I think this application doesn’t have something like that. So I was straightforward and okay, you want it, and then you get it.m — E5
@@
%I mostly wanted to just work with the scheme and the functionality that I had. And the application already had a permission system running and working, so I just kept my solutions to that and hoped that this would already be compliant.m — E7
@@
%Yeah, I did google a lot of stuff to understand how to work with Angular and implement the features. And that might be problematic to privacy compliance because I’m looking. First of all, I don’t know the framework really well. So I’m more inclined to copy solutions, copy code, copy snippets and then not know if that will be a risk.m — C7
@@
%In the first few tasks, I was really confident that it’s data secure. And then the last ones, I had problems implementing the things. So, I think I would have started looking for this if I finished the tasks. Yeah, if I got the functionality right.m — E6
@@
%Because if I’m being completely honest, I don’t know exactly how I should implement this properly in terms of data protection, something like that or what is okay or not okay.m — P8
@@
%That was a really big no-no, I thought. I implemented the function anyway, or at least an attempt at it, and yes, I will say, I find the idea very difficult anyway, but if you really want to have it in, then you would actually have to implement it in such a way that… No, I don’t know how to do this in a data protection-compliant way. In my opinion, the basic idea is somehow gone.m — P8
@@
%But I think it would be in general better if the lawyers knew concepts of IT and new base functionalities of databases and other things like network traffic. So they have a better idea.m — E7
@@
%Also, the communication doesn’t really come up since we don’t really follow all the proper [procedures].m — E8
@@
%So I think it could certainly be promoted a little more, in the sense of: “Hey, there’s someone you could write to.” Well, I’ve had to deal with him now and then, but I don’t think that all of my colleagues have that kind of contact.m — E10
@@
%I would prefer to delegate such tasks directly to other people. I don’t feel like it, I’ll be honest. My head is just way too invested in it, and I don’t want to be prosecuted if there is any small mistake. Then you’re done.m — C4
@@
%Maybe, I think I would have been more hesitant to delete the database, but that doesn’t… I don’t think so.m — E6
@@
%And I wouldn’t even be that surprised if the CTO would have requested to implement a feature like task four so the advertisement emails. I would definitely have said that this is not conforming to the regulation and that this feature can’t be implemented in that way.m — C2
@@
%When it comes to my real life work, if I were to work in an organization, I would expect to be trained or, like, taught exactly what’s sensitive, what’s not sensitivem — C7
@@
%And obviously, the organization itself would put restraints on my account on my profilem — C7
@@
%Because I was too lazy in most of the functions, especially those get all users were already implemented, I just used them straight out of the box. I was thinking about that it might not be the best idea, but I was too lazy at that point.m — C5
@@
%But for how to implement such a feature again, like Stack Overflow or Google, for the privacy part, I don’t knowm — C1
@@
%Like most things, it is very much legalese. If you don’t have the time to find parallel sources, then it’s really difficult to understand what the state actually wants and what it wants to express with it.m — C4
@@
% […] it was a little bit my impression that it would take quite a long time if I worked my way through it now, and that’s maybe what I have right now. There wasn’t enough time left, and that’s why I left it alone. Well, I’ll say, at least at first glance, it’s not really clear what you have to stick to if I’m being honest.m — P8
@@
%Yes, well, maybe a guideline like that, or especially for such, more common data protection use cases, I would say, where you can perhaps look it up briefly, that might be something.m — E10
@@
%What would be, in my opinion, the most helpful would be, for example, for the backup system, some official page or even semi-official page where an example backup system is provided for an abstract technical implementation and the steps that one has to take to make sure that the data is saved in a law-abiding way.m — C2
@@
%Of course, they need it to be backuped, but I would probably encrypt the backup so that the data is safe, of course, but it can be accessed from like, easily accessed from somewhere.m — P5
@@
%I’m more on the other side to make the application safe for hackers and something like that, not data privacym — P7
@@
%I began to implement the function about sending the email advertisements, and I stopped mid-implementation and to just remember what I, what the f**k I’m going to do. This isn’t okay. So it takes some time to get accustomed to such situations.m — P2
@@ 5 Discussion
The field of psychology recognizes expert-layperson communication as challenging [56, 57], and similar issues have been observed in the field of security [58, 59]. For instance, scientific literature often argues for more security advice and more cooperation with security experts (e.g., [60, 61, 62]). However, our study suggested that access to privacy experts has only a small effect on privacy tasks. Several reasons might influence the participants’ behavior in this regard. First, past studies suggested that developers often expressed feeling restricted by privacy requirements in their implementation [5]. Second, privacy experts have been described by our participants as lacking technical knowledge of implementations. This also extends to the privacy requirements participants got from these experts, which have often been described as hard to understand and written in legal language, lacking clear instructions for the developers (see Section 4.5).
@@
In fact, Task~3.2 resulted in the most privacy-compliant solutions, i.e., $6$/30 fully and $3$/30 partially compliant solutions. This might be caused by the option to directly reuse existing functionality for consenting to provide access for doctors to a patient’s health record, which was used by $5$/30 participants. Both Task~3.2 and Task~3.2 required more effort by participants. For the former, they would have to edit existing backups. For the latter, they would have needed to implement a new consent functionality from scratch. This mirrors participants’ interview responses (cf. Section 4.8), where some preferred to reuse existing functionality. Further, they noted that implementing a privacy codebase might be difficult at this stage, claiming it should have been implemented earlier and that they did not consider it part of their tasks, referring to Privacy by Design solutions.
@@ 6 Conclusion
Functionality First, Security Second, and Privacy Third. Many participants showed a functionality-first approach referring to security when asked about privacy issues, with a tendency to adapt a (non)existing level of privacy implementation. However, focusing on potential attackers often does not help user data issues. We argue privacy and security should not be treated as the same issue, as both usually require independent considerations and measures. We propose conducting more research on how to improve privacy-compliant implementation. While Privacy by Design [63] was proposed to be a guiding framework, more research is needed with tools suggesting to support developers with privacy-compliant implementation [64, 65, 66, 67, 68, 69, 70].
@@ Acknowledgments
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2092 CASA - 390781972. % %TC:ignore

### v3

@@ 4.5 Privacy Expert Chat
One participant asked if the backup should be encrypted during Task 3.2. For Task 3.2, two participants asked the expert to clarify if data from the backup also needed to be deleted. For Task 3.2, two participants asked about data minimization, purpose limitation, and access control measures for doctors. One participant contacted the expert with a technical question, as they deleted more data from the backup by accident and were unsure how to resolve the issue. Lastly, concerning Task 3.2, two participants asked questions regarding consent. One participant asked the expert to clarify if the users already had consented to advertisements and to confirm that, without consent, the sending of the emails would not be compliant:
But if a user has not given any consent then sending mails to them would not be privacy-compliant or is it? — E1 Chat
@@
Would you consider it to be okay just to send all the existing users the advertisement email, but include a link to opt-out on future emails of that kind? I am personally still concerned about such an approach, but you are the expert regarding privacy topics - so I’m happy to hear your advice — E9 Chat
@@
I asked the one time, if all the data, if the patient’s data should also be deleted from the backups, but it was the expert, of course, said yes, and it was kind of obvious that it should be deleted — E6
@@
Thanks again and sorry for bugging you so much lately — E9 Chat
@@
It would have been easier for me if either the task was basically in line with data protection or if I had received more precise instructions from the expert — E9
@@
Cause I’m kind of afraid of asking too much — E2
@@
That’s funny. Yeah. Maybe I was just too lazy and wanted to get the tasks done — E4
@@
I don’t know because I have no… Yeah, I should have used it. Yeah, I think I should have used it. — E2
@@ 4.6 Confidence
Twenty participants found the programming tasks moderately challenging. Two (P6, E5) found them not challenging at all, and three (C1, C3, E3) found them slightly challenging. Four (C7, P5, E1, E6) found the tasks very challenging, and for one participant (C4), the tasks were extremely challenging. An overview of the self-evaluation for each participant can be found in Table IV.
@@ 4.8 Participants’ Explanation for Privacy Behavior
I just found something, but I haven’t written it because it’s already in the project, so it’s not my fault — E2
indicating that participants while noticing privacy issues, decided not to address them as they were not seen as part of a developer’s responsibility. For example, one participant commented on Task 3.2:
I’m not sure about whether it’s okay to send everyone an email just randomly. So I don’t know whether you had to make some subscription to something like a newsletter or something because, yeah, I don’t know. And I think this application doesn’t have something like that. So I was straightforward and okay, you want it, and then you get it. — E5
@@
I mostly wanted to just work with the scheme and the functionality that I had. And the application already had a permission system running and working, so I just kept my solutions to that and hoped that this would already be compliant. — E7
@@
Yeah, I did google a lot of stuff to understand how to work with Angular and implement the features. And that might be problematic to privacy compliance because I’m looking. First of all, I don’t know the framework really well. So I’m more inclined to copy solutions, copy code, copy snippets and then not know if that will be a risk. — C7
@@
In the first few tasks, I was really confident that it’s data secure. And then the last ones, I had problems implementing the things. So, I think I would have started looking for this if I finished the tasks. Yeah, if I got the functionality right. — E6
@@
Because if I’m being completely honest, I don’t know exactly how I should implement this properly in terms of data protection, something like that or what is okay or not okay. — P8
@@
That was a really big no-no, I thought. I implemented the function anyway, or at least an attempt at it, and yes, I will say, I find the idea very difficult anyway, but if you really want to have it in, then you would actually have to implement it in such a way that… No, I don’t know how to do this in a data protection-compliant way. In my opinion, the basic idea is somehow gone. — P8
@@
But I think it would be in general better if the lawyers knew concepts of IT and new base functionalities of databases and other things like network traffic. So they have a better idea. — E7
@@
Also, the communication doesn’t really come up since we don’t really follow all the proper [procedures]. — E8
@@
So I think it could certainly be promoted a little more, in the sense of: “Hey, there’s someone you could write to.” Well, I’ve had to deal with him now and then, but I don’t think that all of my colleagues have that kind of contact. — E10
@@
I would prefer to delegate such tasks directly to other people. I don’t feel like it, I’ll be honest. My head is just way too invested in it, and I don’t want to be prosecuted if there is any small mistake. Then you’re done. — C4
@@
Maybe, I think I would have been more hesitant to delete the database, but that doesn’t… I don’t think so. — E6
@@
And I wouldn’t even be that surprised if the CTO would have requested to implement a feature like task four so the advertisement emails. I would definitely have said that this is not conforming to the regulation and that this feature can’t be implemented in that way. — C2
@@
When it comes to my real life work, if I were to work in an organization, I would expect to be trained or, like, taught exactly what’s sensitive, what’s not sensitive — C7
@@
And obviously, the organization itself would put restraints on my account on my profile — C7
@@
Because I was too lazy in most of the functions, especially those get all users were already implemented, I just used them straight out of the box. I was thinking about that it might not be the best idea, but I was too lazy at that point. — C5
@@
But for how to implement such a feature again, like Stack Overflow or Google, for the privacy part, I don’t know — C1
@@
Like most things, it is very much legalese. If you don’t have the time to find parallel sources, then it’s really difficult to understand what the state actually wants and what it wants to express with it. — C4
@@
[…] it was a little bit my impression that it would take quite a long time if I worked my way through it now, and that’s maybe what I have right now. There wasn’t enough time left, and that’s why I left it alone. Well, I’ll say, at least at first glance, it’s not really clear what you have to stick to if I’m being honest. — P8
@@
Yes, well, maybe a guideline like that, or especially for such, more common data protection use cases, I would say, where you can perhaps look it up briefly, that might be something. — E10
@@
What would be, in my opinion, the most helpful would be, for example, for the backup system, some official page or even semi-official page where an example backup system is provided for an abstract technical implementation and the steps that one has to take to make sure that the data is saved in a law-abiding way. — C2
@@
Of course, they need it to be backuped, but I would probably encrypt the backup so that the data is safe, of course, but it can be accessed from like, easily accessed from somewhere. — P5
@@
I’m more on the other side to make the application safe for hackers and something like that, not data privacy — P7
@@
I began to implement the function about sending the email advertisements, and I stopped mid-implementation and to just remember what I, what the f**k I’m going to do. This isn’t okay. So it takes some time to get accustomed to such situations. — P2
@@ 5 Discussion
The field of psychology recognizes expert-layperson communication as challenging [bromme2000beyond, bromme2005barriers], and similar issues have been observed in the field of security [poller2017can, thomas2018security]. For instance, scientific literature often argues for more security advice and more cooperation with security experts (e.g., [gutfleisch2022does, weir2020from, busse2019replication]). However, our study suggested that access to privacy experts has only a small effect on privacy tasks. Several reasons might influence the participants’ behavior in this regard. First, past studies suggested that developers often expressed feeling restricted by privacy requirements in their implementation [horstmann2024those]. Second, privacy experts have been described by our participants as lacking technical knowledge of implementations. This also extends to the privacy requirements participants got from these experts, which have often been described as hard to understand and written in legal language, lacking clear instructions for the developers (see Section 4.5).
@@
In fact, Task 3.2 resulted in the most privacy-compliant solutions, i.e., 66/30 fully and 33/30 partially compliant solutions. This might be caused by the option to directly reuse existing functionality for consenting to provide access for doctors to a patient’s health record, which was used by 55/30 participants. Both Task 3.2 and Task 3.2 required more effort by participants. For the former, they would have to edit existing backups. For the latter, they would have needed to implement a new consent functionality from scratch. This mirrors participants’ interview responses (cf. Section 4.8), where some preferred to reuse existing functionality. Further, they noted that implementing a privacy codebase might be difficult at this stage, claiming it should have been implemented earlier and that they did not consider it part of their tasks, referring to Privacy by Design solutions.
@@ 6 Conclusion
Functionality First, Security Second, and Privacy Third. Many participants showed a functionality-first approach referring to security when asked about privacy issues, with a tendency to adapt a (non)existing level of privacy implementation. However, focusing on potential attackers often does not help user data issues. We argue privacy and security should not be treated as the same issue, as both usually require independent considerations and measures. We propose conducting more research on how to improve privacy-compliant implementation. While Privacy by Design [cavoukian2009privacy] was proposed to be a guiding framework, more research is needed with tools suggesting to support developers with privacy-compliant implementation [AndMahWhiEnc+20, SheSuKanPha+23, li2018coconut, li2021honeysuckle, FerBriTia+23, KleRolBarKar+23, SrdBasHub23].
@@ Acknowledgments
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2092 CASA - 390781972.